OpenStudy (anonymous):

f(x)=6^x, what's the formular when shifting f(x) 2 units to the right, thank you

OpenStudy (anonymous):

f(x)=6^(x-2)

OpenStudy (anonymous):

thank you, so smart, what about reflecting f(x)=6^x about the x-axis

OpenStudy (anonymous):

f(x) = -6^x

OpenStudy (anonymous):

thank you, you are amazing!

OpenStudy (anonymous):

i know that.. :D